OpenStudy (anonymous):

Solve the system by elimination- {-2x+2y+3z=0 {-2x-y+z=-3 {2x+3y+3z=5 Can somebody answer this step by step or help me figure it out?

OpenStudy (anonymous):

Take the first two equations.... -2x + 2y + 3z = 0 -2x - y + z = -3 -->(-1)-2x - y + z = - 3 ---------------- -2x + 2y + 3z = 0 2x + y - z = 3 (result of multiplying by -1) ----------------add 0 + 3y + 2z = 3 3y + 2z = 3 now take the last two equations (you have to eliminate the x's) -2x - y + z = - 3 2x + 3y + 3z = 5 ----------------add 0 + 2y + 4z = 2 2y + 4z = 2 now take the answers you got an eliminate a variable 3y + 2z = 3 -->(-2)3y + 2z = 3 2y + 4z = 2 ----------- -6y - 4z = - 6 (result of multiplying by -2) 2y + 4z = 2 -----------add -4y = - 4 -4y/-4 = -4/-4 y = 1 now sub 1 in for y in either of the equations with two variables 3y + 2z = 3 3(1) + 2z = 3 3 + 2z = 3 2z = 3 - 3 2z = 0 z = 0/2 = 0 now sub known variables into any of the original equations 2x + 3y + 3z = 5 2x + 3(1) + 3(0) = 5 2x + 3 + 0 = 5 2x = 5 - 3 2x = 2 x = 1 now lets check... -2x + 2y + 3z = 0 -2(1) + 2(1) + 3(0) = 0 -2 + 2 + 0 = 0 0 = 0 (correct) lets double check... -2x - y + z = -3 -2(1) - 1 + 0 = - 3 -2 - 1 + 0 = - 3 -3 = - 3 (correct) x = 1, y = 1, z = 0......(1,1,0) Do you have any questions ?

OpenStudy (anonymous):

Thank you sooo much...too bad I gave up and skipped the question :/ but thank you!

OpenStudy (anonymous):

its not a problem........its not that hard, just time consuming....lol